OpenStudy (anonymous):

help please From the top of a 175m hill the angle of depression of a house is 75. From the same point on the hill the angle of depression to a church is 62. Assuming the house and church are horizontally in a straight line on the same side of the hill, what is the distance from the house to the church?

OpenStudy (anonymous):

distance to house = 175 / (tan 75) = 175/3.732 = 46.9 M distance to church = 175 / (tan 62) = 175/1.881 =93.05M distance from house to church = 93.05-46.9 = about 46m

OpenStudy (anonymous):

can you help me draw the diagram

OpenStudy (anonymous):

 sorry, don't know how to make the drawing work

OpenStudy (anonymous):

but in what I was able to draw, the bottom of the triangle can be found if the height (175) and the angle (62 or 75) is known... i.e., tan (angle) = height / ground distance - or - ground distance = height / tan(angle)